### v1

# A Systematic Investigation of Knowledge Retrieval and Selection for Retrieval Augmented Generation

Abstract
Retrieval-augmented generation (RAG) has emerged as a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. In this paper, we perform a comprehensive analysis of how knowledge retrieval and selection influence downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge retrieval and selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing a limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
A Systematic Investigation of Knowledge Retrieval and Selection for Retrieval Augmented Generation
Xiangci Li1, 2    Jessica Ouyang1 1 University of Texas at Dallas, 2 Amazon Web Services lixiangci8@gmail.com, jessica.ouyang@utdallas.edu
**footnotetext:  Work performed outside of AWS.
1 Introduction
Retrieval-augmented generation (RAG) has become a pivotal technique in natural language generation, enhancing a language model’s ability to produce relevant, informed output by incorporating external knowledge Gao et al. (2023); Fan et al. (2024). This approach complements the model’s internal knowledge, which is inherently constrained by the information that was available during training, by retrieving up-to-date, relevant information to support and inform its output.
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2). We hypothesize that knowledge selectors may not always improve downstream generation performance, and that there may be a selection bias where only positive results involving knowledge selector modules are published, while experiments where knowledge selectors are not helpful simply do not report those results.
In this work, we perform a systematic analysis of how knowledge retrieval and selection affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different retrieval and selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
Figure 1: Our simulation experiment pipeline.
2 Related Work
Retrieval-augmented generation (RAG)
has been extensively studied in recent years. Early works Guu et al. (2020); Lewis et al. (2020b); Shuster et al. (2021) jointly fine-tuned a dense retriever (e.g. DPR Karpukhin et al. (2020)) and generator (e.g. BART Lewis et al. (2020a)), which required dedicated training datasets. More recently, the introduction of large language models (LLMs) has made RAG more convenient to implement due to their strong generation performance, in-context learning ability Brown (2020); Kojima et al. (2022), and drastically longer context windows — from BART’s 1024 tokens to more than a million for modern LLMs Lee et al. (2024). As a result, recent RAG research has shifted to using LLMs Gao et al. (2023); Fan et al. (2024); in this work, we follow this trend to focus on LLM-based RAG.
Knowledge-grounded dialogue generation
Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, despite these works demonstrating the advantages of their overall approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based RAG works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To the best of our knowledge, there is no prior work explaining this gap in knowledge selection. The closest work to ours is Cuconasu et al. (2024) which studies the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via simulations, rather than being restricted to specific retriever or selector implementations.
3 Approach
3.1 Task Formulation
RAG involves three steps: (1) Knowledge retrieval, where a retriever module retrieves a set of candidate knowledge $K$ based on a query $q$. This step aims to retrieve as much knowledge that is relevant to the query as possible, attempting to balance knowledge recall and precision. (2) Optionally, knowledge selection removes retrieved knowledge that is less relevant to further improve the knowledge precision, producing $K^{\prime}\subseteq K$. (3) Finally in the generation step, the generator takes the query $q$ and the selected retrieved knowledge $K^{\prime}$ to generate the output text $r$.
The knowledge in $K$ and $K^{\prime}$ can be heterogeneous and come from multiple sources, such as external documents, knowledge graphs, or conversation histories. This RAG framework can be applied to various tasks, such as dialogue generation, question answering, fact verification, and code generation Gao et al. (2023); Fan et al. (2024). In our experiments on dialogue generation and question answering, we simulate steps 1 and 2 by creating controlled knowledge sets $K^{\prime}$ and observing the resulting performance of step 3.
3.2 Knowledge Simulation
We aim to perform a systematic analysis of the effect of knowledge retrieval and selection outcomes on the downstream RAG performance by simulation. As Figure 1 shows, for each query, given a fixed pool of available knowledge with gold knowledge annotations, we sample gold knowledge and distractor knowledge at varying rates $p_{gold}$ and $p_{noise}$ to precisely simulate a wide range of quality for the retrieved and selected knowledge $K^{\prime}$, from noise only to gold knowledge only. For example, if $p_{gold}=0.5$, then each piece of gold knowledge has a 50% chance to be sampled. In this way, we simulate a wide distribution of knowledge retrievers and selectors of different performance levels, as measured by knowledge precision and recall, which is computed based on the annotated gold knowledge. We then test the performance of the generator model given the different quality of retrieved and selected knowledge $K^{\prime}$.
4 Experimental Settings
4.1 Datasets
While there are several datasets supporting the RAG framework, relatively few provide high-quality, human-annotated gold knowledge. We use two popular and representative datasets, Wizard of Wikipedia (WoW; Dinan et al., 2019) and HotpotQA Yang et al. (2018) for our experiments.
WoW
is an open-domain dialogue dataset based on Wikipedia knowledge that has been widely used in prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023). The wizard speaker (a human annotator playing the part of the dialogue system) has access to Wikipedia knowledge, while the apprentice speaker does not. The apprentice is given a starting conversation topic, but otherwise speaks freely. The last two turns of dialogue are used as the query to retrieve Wikipedia passages ($K$), and the wizard selects a single knowledge sentence ($K^{\prime}$) to generate their response.
Despite its popularity, this dialogue generation task is challenging to evaluate. First, because the wizard is limited to selecting only a single knowledge sentence, it is possible for the unselected “distractor" knowledge to also be relevant to the wizard’s response. Second, as is the case in many natural language generation tasks, the gold responses in WoW are not the only plausible responses, making it harder to quantify the correctness of a generated response. Nonetheless, WoW is a well-annotated dialogue dataset that is close to a real-life scenario where the gold knowledge and responses are noisy.
HotpotQA
is a question-answering dataset derived from Wikipedia knowledge that we include in our experiments to mitigate the ambiguity of WoW. The multi-hop questions and answers are first derived from gold knowledge graphs, and then distractor knowledge is injected to create noise. As a result, unlike WoW, the answers in HotpotQA are strongly dependent on the gold knowledge, and the distractor knowledge is unlikely to be relevant. The short and unambiguous gold answers make it simpler to evaluate via F1 scores.
4.2 Generators
To align with the latest trends in RAG research, as well as to make the analysis simple and generally replicable, we adopt LLMs as our generators. We use three API-based LLMs, OpenAI GPT-4o-mini, LLaMA 3.1 8B, and Mistral 7B-Instruct, as our generator models; the varying sizes of the LLMs allow us to investigate the impact of generator complexity on knowledge usage.
Input Knowledge KP KR KF1 R-L F1
GPT-4o-mini
No knowledge 0 0 0 0.110 0.200
Full knowledge 0.015 1 0.031 0.140 0.251
Gold knowledge 1 1 1 0.167 0.276
LLM selector 0.054 0.720 0.101 0.139 0.247
LLaMA 3.1 8B
No knowledge 0 0 0 0.111 0.216
Full knowledge 0.015 1 0.031 0.138 0.248
Gold knowledge 1 1 1 0.164 0.278
LLM selector 0.053 0.317 0.009 0.126 0.231
Mistral 7B Instruct
No knowledge 0 0 0 0.113 0.203
Full knowledge 0. 1 0. 0.131 0.233
Gold knowledge 1 1 1 0.172 0.268
LLM selector 0.036 0.522 0.067 0.123 0.221
Table 1: WoW response generation performance benchmarked by different LLM generatorss. We measure knowledge precision (KP), recall (KR), and F1 (KF1); response ROUGE-L F1 (R-L); and response F1.
Input Knowledge KP KR KF1 EM F1
GPT-4o-mini
No knowledge 0 0 0 0.330 0.437
Full knowledge 0.059 1 0.111 0.668 0.780
Gold knowledge 1 1 1 0.710 0.828
LLM selector 0.351 0.685 0.464 0.622 0.724
LLaMA 3.1 8B
No knowledge 0 0 0 0.200 0.298
Full knowledge 0.059 1 0.111 0.372 0.545
Gold knowledge 1 1 1 0.414 0.671
LLM selector 0.304 0.405 0.347 0.294 0.469
Mistral 7B Instruct
No knowledge 0 0 0 0.208 0.260
Full knowledge 0.059 1 0.046 0.668 0.151
Gold knowledge 1 1 1 0.502 0.627
LLM selector 0.06 0.244 0.103 0.186 0.245
Table 2: HotpotQA answer generation performance benchmarked by different LLM generators. We measure knowledge precision (KP), recall (KR), and F1 (KF1); answer exact match (EM); and answer F1.
4.3 Knowledge Sampling
Since knowledge precision and recall are the most common metrics for knowledge retrieval and selection performance, we use these metrics as the basis for our analysis. We find that $p_{gold}$ linearly correlates with knowledge recall scores, while $p_{noise}$ exponentially correlates with knowledge precision. Thus, to simulate retrievers/selectors with specific knowledge precision and recall scores, we use grid search in the linear space of $p_{gold}$ and both the linear and exponential spaces of $p_{noise}$ to ensure most grids in the knowledge precision-recall space are covered by our experiments.
As Figure 1 shows, both WoW and HotpotQA provide a retrieved knowledge set $K$ for each query $q$ such that the gold knowledge is a subset of $K$. In both datasets, $K$ is retrieved using bigram TF*IDF Chen et al. (2017) and contains text passages, where each passage consists of an article title and a few knowledge sentences. We perform knowledge sampling at the sentence level; the sampled knowledge simulates the end result of applying both the knowledge retrieval and selection steps.
In our experiments, we refer to using the entire provided knowledge set $K$ as the input knowledge to the generator (i.e. no knowledge selection) as the “full knowledge” setting. We also include the “no knowledge" setting, where the generators receive no external knowledge at all, as a weak baseline, and “gold knowledge," corresponding to perfect knowledge selection, as an upper bound.
Figure 2: Scatter plot of response/answer F1, plotted against knowledge precision (x-axis) and recall (y-axis), by GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom). The left column shows results on WoW; the right shows HotpotQA. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting, while those highlighted in cyan indicate settings underperforming the “no knowledge" setting.
4.4 Implementation Details
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.1) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18***https://platform.openai.com/docs/models/gpt-4o-mini, Together AI***https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW.
5 Simulation Results
Overall we observe several consistent trends, except for Mistral-7B-Instruct, which is a relatively weak-performing LLM.
RAG for LLM is beneficial.
As Tables 1 & 2 show, generators without any retrieved knowledge (“no knowledge" setting) perform poorly on WoW and HotpotQA, even though they may have pre-trained on the Wikipedia articles that WoW and HotpotQA are derived from. This finding indicates the LLMs are not over-fitted on the WoW and HotpotQA datasets, and applying RAG is beneficial.
Interestingly, our HotpotQA results in Figure 2 show that distractor knowledge significantly harms performance; the cyan dots show that generators receiving mostly distractor knowledge underperform the “no knowledge” setting, which uses only the LLM’s internal knowledge. This trend is not observed with WoW, supporting our observation that HotpotQA’s distractor knowledge is truly irrelevant, while the “distractor" knowledge in WoW may only be less relevant than the single gold sentence, but not completely irrelevant.
Figure 3: Scatter plot of HotpotQA answer F1 versus knowledge precision (top), knowledge recall (middle), and knowledge F1 (bottom). The left column shows GPT-4o-mini as the generator; the right column shows Mistral-7B-Instruct. Plots for LLaMa 3.1 8B and the WoW dataset are in Appendix A.
Full knowledge setting without knowledge selection is a strong baseline.
Tables 1 & 2 show that the “full knowledge" setting, which corresponds to knowledge retrieval with perfect recall, but no knowledge selection, is a very strong baseline. The tables also include an LLM knowledge selector, implemented with the same LLM as the generator (prompts in Appendix A.1), to show the performance of a typical knowledge selector module. None of these LLM knowledge selectors outperform the “full knowledge” setting, except Mistral-7B-Instruct on HotpotQA.
This finding mostly contrasts with those of prior knowledge selection works Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024). For example, in Table 2, we find that for GPT-4o-mini, a very strong generator model, the “full knowledge" setting achieves 0.780 answer F1 on HotpotQA, only 0.048 lower than the “gold knowledge" F1; similar observations are seen for LLaMA 3.1 8B and for both models on WoW. Since the performance gap between the “full knowledge” and “gold knowledge” settings is the space for a knowledge selector to improve generation performance, we conclude that strong generators simply have less room for improvement via knowledge selection.
Knowledge precision & recall together are good indicators of generation performance.
Figure 2 shows that generation performance varies smoothly with knowledge precision and recall, indicating that knowledge precision and recall together are major determinants of generation performance. Due to the pipeline nature of the retriever, selector, and generator components of RAG, knowledge recall can only be improved by the retriever. The selector can only improve knowledge precision, but it is likely to also reduce recall. Therefore, we can visualize the effect of applying a knowledge selector: it moves the RAG system towards the lower-right corner of the plots in Figure 2; the better the knowledge selector, the more it will move right, and the less it will move down.
Knowledge recall is the most crucial knowledge metric for strong generators.
We find that for strong generator models, the knowledge recall score is the best single knowledge metric for estimating generation performance; Figure 3 shows a very strong correlation between knowledge recall and answer F1 for GPT-4o-mini on HotpotQA, and Figure 8 in the Appendix shows the same for LLaMA 3.1 8B and both models on WoW. Moreover, knowledge recall is the most important factor in improving generation performance given a fixed generator, while the knowledge precision score is a secondary factor.
Figure 4 further shows that increases in knowledge recall correspond to significant increases in answer F1 scores. Taking GPT-4o-mini on HotpotQA as an example, we see that moving from the left end of a color contour to the right (i.e. keeping knowledge recall fixed while improving precision) only slightly improves answer F1. In contrast, moving from one contour to another (i.e. varying knowledge recall) significantly impacts answer F1 for non-zero precision scores. Thus, to improve knowledge quality for RAG, improving the retriever’s recall score is the top priority; improving precision via a knowledge selector has a limited contribution to generation performance, especially for strong generators.
For weaker generators, like Mistral 7B-Instruct, the relationship between generation and knowledge F1 is stronger, while the correlation with recall is weaker (Figure 3, right column). Similarly, we see much less separation between color contours (recall) and a much steeper increase in answer F1 with precision (Figure 4, bottom row).
Figure 4: Color contours of answer F1 versus the knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA. Each contour represents a different knowledge recall score; moving left to right visualizes improving the performance (precision) of the knowledge selector.
Generator capability determines both overall performance and the usefulness of knowledge selection.
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards***https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input knowledge and thus requires a selector to filter out distractor knowledge — knowledge F1 best correlates to the answer F1 for Mistral-7B-Instruct in Figure 3.
Task and dataset are also key factors in RAG performance.
Figures 2 & 4 show that the same generator can show drastically different performance trends between WoW and HotpotQA. For example, while Mistral-7B-Instruct’s performance degrades without a knowledge selector on HotpotQA, this is not the case for WoW.
In addition, Figure 4 shows that attempting to improve the knowledge selector on top of a weak retriever (i.e. increasing precision when recall is low) hurts generation performance on WoW, as well as for GPT-4o-mini on Hotpot QA. In this scenario, more total knowledge, regardless of noise, improves response generation. These counter-intuitive observations are likely due to the nature of the task and the annotation quality. The solution space in HotpotQA is small given a specific question, and HotpotQA has a cleaner separation between gold and distractor knowledge, whereas in WoW, there are more plausible responses given the same conversation history and knowledge, as well as “distractor" knowledge that may actually be relevant to the gold response.
Non-monotonic trends in improving the knowledge selector.
Interestingly, we observe that the boundary of where the knowledge selector improves generation performance (the border between the orange and white areas in Figure 2) is convex for all three LLMs on WoW. This phenomenon can also be seen in Figure 4, where some color contours intersect with the “full knowledge answer F1” baseline (dashed red line) multiple times. In other words, we observe that generation performance is non-monotonic as knowledge precision increases with fixed knowledge recall. The fact that this phenomenon is only observed on WoW may be due to the relatively noisy gold knowledge annotations in WoW; we can produce similar behavior by artificially injecting noise into HotpotQA’s gold knowledge annotations (Appendix A.3).
6 Conclusion and Discussion
In this study, we systematically examined the behavior of RAG generation in relation to the quality of retrieved knowledge.
6.1 Overall Observations
Summing up the results in Section 5, we find an interaction effect between generator capability and the ambiguity of the task/dataset on the RAG generation performance. We find two types of generator behavior:
A strong generator model can achieve good performance without a knowledge selector because it is robust to noise, and thus performs better the more gold knowledge is retrieved, despite the presence of distractor knowledge. Thus, knowledge recall correlates well with generation F1, and there is not much room for improvement through the addition of a knowledge selector.
A weak generator model cannot handle distractor knowledge and requires a selector to refine the noisy input knowledge, and thus knowledge F1 correlates well with generation F1. Since nowadays most popular generator models are strong LLMs, knowledge selectors have a limited benefit. We hypothesize that prior work that found performance improvements from dedicated knowledge selectors saw those benefits because their generator models were weak; most prior work used BART Lewis et al. (2020a) as the generator.
Finally, we note that the strength of a generator model is relative; even a SOTA generator can fall into the “weak" category, given a sufficiently noisy and challenging task/dataset. We hope the visualizations shown in our figures can help guide future practitioners improve their RAG systems in real-world applications.
6.2 Recommendations to Future Practitioners
Based on our observations, we make the following recommendations for future practitioners considering how to improve performance in a real-world RAG scenario.
First, benchmark the generation performance without external knowledge, with all candidate knowledge, and if possible, with gold knowledge only. The “no knowledge” setting measures the generator’s base performance, “full knowledge” serves as a strong baseline corresponding to knowledge retrieval with perfect recall, and “gold knowledge” gives the upper bound of RAG performance with perfect knowledge selection. The gap between “full knowledge" and “gold knowledge" is the potential performance gain brought by adding a knowledge selector.
To improve RAG performance, increasing the knowledge recall score is the most effective strategy. Thus, for modern LLM generators, improving the knowledge retriever’s recall is key. In practice, the retriever or selector may encounter false-negative gold knowledge during training, as is the case with WoW, which makes prioritizing recall even more important. Moreover, because modern LLM generators have long maximum input windows (e.g. 128k for GPT-4o & -mini), having too much knowledge is less likely to be a problem.
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
Limitations
Computational resources for simulations.
Due to the high cost of larger-scale simulation experiments using API-based LLMs, we use only a subset of the WoW and HotpotQA for experiments. As a result, our data subset may contain some minor noise during knowledge sampling and LLM generation, and the contours in Figure 4 are not smooth. However, such noise is not likely to affect our conclusions. For a similar reason, we only chose three LLMs for our experiments. As a result, we may have missed out on more subtle phenomena that can only be seen from results on a larger number of generators.
Datasets for in-depth analysis.
In addition to the cost issue, there are very few RAG datasets with human-annotated knowledge Friel et al. (2024), which further limits our simulation experiment settings. Furthermore, even though we regard WoW as a relatively noisy dataset in this work compared to HotpotQA, to the best of our knowledge, WoW is one of the most cleanly annotated datasets among all datasets. Therefore, given our hypothesis in Section 5 that WoW has a larger solution space than HotpotQA, we do not have numerical evidence because each example in WoW only contains one gold response thus we cannot measure the size of the solution space.
Ethics Statement
Since all datasets and LLMs used in this work are publicly available or accessible, and no data collection introducing sensitive information is performed, the ethical considerations of this study are minimal.
Appendix A Appendix
A.1 Prompts
Tables 3 & 4 show the prompts we use for response generation on WoW and HotpotQA, respectively. Tables 5 & 6 show the prompts we use for the “LLM selector" knowledge selection baseline on WoW and HotpotQA, respectively. While we experimented with Chain-of-Thought Wei et al. (2022) for these prompts, they did not outperform zero-shot prompting for LLaMA 3.1 8B and Mistral-7B-Instruct, so we use zero-shot prompting throughout our experiments to keep the settings as simple as possible.
A.2 Additional Plots
Figures 7, 8, and 9 extend Figure 3, with knowledge precision, knowledge recall, and knowledge F1 plotted against response/answer F1 for WoW and HotpotQA, respectively, for all three LLM generators. Figure 8 shows that knowledge recall correlates strongly with response/answer F1 for both GPT-4o-mini and LLaMA 3.1 8B on both WoW and HotpotQA, as well as for Mistral 7B-Instruct on WoW, while Figure 9 shows that knowledge F1 has the stronger correlation for Mistral 7B-Instruct on HotpotQA.
A.3 Noisy HotpotQA
As we discuss in Section 4.1, the WoW gold knowledge annotations are noisy in that only one sentence can be labeled as gold knowledge, even if other sentences are still relevant. To test the impact of such knowledge annotation noise on our RAG simulation results, we deliberately inject noise into HotpotQA: we drop one sentence from the gold knowledge sentences for each example to create a noisy HotpotQA dataset. As a result, the average number of gold sentences per example drops from 2.4 to 1.4, and a small portion of the “distractor" knowledge in the noisy dataset is actually gold knowledge from the original HotpotQA (i.e. false-negative knowledge).
As Figures 5 & 6 show, the injected noise drastically impairs the knowledge selector’s efficacy: even given a “perfect” knowledge selector with a knowledge precision score of 100%, the answer F1 still underperforms the “full knowledge” setting. This is because our knowledge precision and recall scores are computed based on the labeled gold knowledge sentences, so if the labels are noisy, the knowledge metrics are corrupted while the answer F1 is intact. Note that Figure 6 presents a non-monotonic trend in improving knowledge selector, just as we observed in Section 5 for WoW, verifying that noisy knowledge annotations cause this non-monotonic behavior.
We argue that the simulation results on WoW and noisy HotpotQA are more reflective of real-world RAG applications than the original (clean) HotpotQA. In a real application, there is no way to know the test-time knowledge recall and precision, and it is likely to be slightly different than the training-time knowledge recall and precision.
Figure 5: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini on noisy HotpotQA.
Figure 6: Color contours of answer F1 versus the knowledge precision for GPT-4o-mini on noisy HotpotQA. Each contour represents a different knowledge recall score; moving left to right visualizes improving the performance (precision) of the knowledge selector.
Prompt
The following is the conversation between the "Wizard", a knowledgable speaker who can access to Wikipedia knowledge sentences to chat to with the "Apprentice", who does not have access to Wikipedia.
The conversation is about “{{persona}}”.
{if history %}
Here is the conversation history:
{for turn in history %}
{{turn.speaker}}: {{turn.text}}
{endfor %}
{endif %}
{if context %}
Here are some retrieved Wikipedia knowledge for the Wizard. The Wizard can choose any subset of the following knowledge. It’s also allowed to not choosing any of them.
{for evidence in context %}
Title: {{ evidence.title }}
Sentences:
{for sentence in evidence.sentences %}
- {{ sentence }}
{endfor %}
{endfor %}
{endif %}
Given the knowledge above, make a very brief, such as one sentence, natural response for the Wizard.
Not all information in the chosen knowledge has to be used in the response. Do not include the speaker’s name in the response.
The Wizard’s response is:
Table 3: Jinja2 prompt template for Wizard of Wikipedia.
Prompt
Answer this question from HotpotQA with a response that is as short as possible, e.g. one word:
{{ question }}
{if context %}
Use the following support evidence to answer:
{for evidence in context %}
Title: {{ evidence.title }}
Sentences:
{for sentence in evidence.sentences %}
- sentence
{endfor %}
{endfor %}
{endif %}
Table 4: Jinja2 prompt template for HotpotQA.
Prompt
The following is the conversation between the "Wizard", a knowledgable speaker who can access to Wikipedia knowledge sentences to chat to with the "Apprentice", who does not have access to Wikipedia.
The conversation is about “{{persona}}”.
{if history %}
Here is the conversation history:
{for turn in history %}
{{turn.speaker}}: {{turn.text}}
{endfor %}
{endif %}
{if context %}
Here are some retrieved Wikipedia knowledge for the Wizard. The Wizard can choose any subset of the following knowledge. It’s also allowed to not choosing any of them.
Title: {{ context.title }}
Sentences:
{for index, sentence in context.sentences | enumerate %}
- {{ index + 1}}. {{ sentence }}
{endfor %}
Answer the decision of selecting each sentence with a list of boolean values. Make sure the length of the list is the same as the number of sentences.
Answer in JSON format so that I can use Python eval() to parse it, e.g.,:
“answer”: [True, False, …, True, False],
Your answer:
Table 5: Jinja2 prompt template for Wizard of Wikipedia knowledge selector.
Prompt
Given this question from HotpotQA:
{{ question }}
For each of the following sentences in the context, judge whether they are relevant to answering the question.
Title: {{ context.title }}
Sentences:
{for index, sentence in context.sentences | enumerate %}
- {{ index + 1}}. {{ sentence }}
{endfor %}
Answer the decision of selecting each sentence with a list of boolean values. Make sure the length of the list is the same as the number of sentences.
Answer in JSON format so that I can use Python eval() to parse it, e.g.,:
“answer”: [True, False, …, True, False],
Your answer:
Table 6: Jinja2 prompt template for HotpotQA knowledge selector.
Figure 7: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.
Figure 8: Scatter plot of answer F1 versus the knowledge recall for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.
Figure 9: Scatter plot of answer F1 versus the knowledge F1 for GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom); the left column shows results on WoW, and the right shows HotpotQA.